OpenStudy (helpmeplease!!!):

The area of a square is 40 cm2. Its side length is enlarged by a scale factor of 3. What is the area of the new square? Area of enlarged square = ____cm2

OpenStudy (anonymous):

multiply by square of scale factor

OpenStudy (helpmeplease!!!):

im new to the topic so i dont really understand

OpenStudy (anonymous):

Area of square : (side)^2 Now, your side has become : 3*Side.. So, Area of square = (3side)^2..

OpenStudy (helpmeplease!!!):

is it just 40 x 3? which is 120 cm^2?

OpenStudy (anonymous):

let us suppose length=2cm area=2^2=4 scale factor=3 new length=2*3=6 cm area=6^2=36 36=4*9 or36=4*3^2

OpenStudy (helpmeplease!!!):

hmm sorry im still not sure

OpenStudy (anonymous):

new area=36 old area=4 scale factor=3 so new area =old area*(scale factor)^2

OpenStudy (skullpatrol):